# Explainable Artificial Intelligence techniques for interpretation of food datasets: a review

### v1

@@ 5.4 Mixed Explanation
Conversely, using spectral data, [153] investigated 1H NMR spectra to determine the geographical origins of Asian red pepper powders, employing ML, SVM, and CNN models with dimensionality reduction techniques. Grad-CAM and SHAP provided insights into the decision-making processes, highlighting metabolite distribution variations as key classification factors. This study demonstrated the potential of these models for broader applications in food authenticity verification.
@@ 6 Explaining Nutritional Value
Table 3: Summary of the works introducing applying XAI for nutritional value surveyed in Section 6, according to their data type and explanation type (labelled as “Expl. type”).
@@
6.1 Visual Explanation
@@
6.2 Numerical Explanation
Spectral data was explored in [175]. The authors employed visible NIR point spectroscopy to estimate sugar content in grape varieties at different maturity stages. Regression ML algorithms and a CNN were applied, with XAI techniques such as Variable Importance in Projection and Gini Importance validating the models and identifying key spectral features. On the other hand, tabular data, was discussed in [30], [176], [32], and [178] apply ML techniques to various food-related challenges. [30] used XGBoost to estimate added sugar content in foods, with SHAP enhancing model transparency for consumer awareness in regions without mandatory labeling. [176] developed the Flavonoid Astringency Prediction Database, employing ML models like RF to explore the relationship between molecular structures and flavor properties. Similarly, [32] applied ML to differentiate pepper spices during storage, using SHAP to identify key organic compounds. [178] developed an XGBoost-based model for predicting drug-food interactions using molecular fingerprint similarities, with SHAP providing insights into influential features relevant to clinical applications and dietary planning. [177] proposed a graph-based ML approach to predict the outcomes of formulation trials, aiming to reduce laboratory experiments, material waste, and development time in food design. To enhance interpretability, they applied GNNExplainer [180], a global explanation method tailored for graph neural networks.
6.3 Rule-based Explanation
Only two significant studies employed XAI techniques to generate rule-based explanations in the context of Nutritional Values. [31] exploited pictorial data to propose a similarity score based on user community preferences, enhancing recommendation quality. The rule-based explainability method assigned each image to an appropriate food diet based on user profiles, supporting personalized dietary recommendations. [179] presented a novel no-code methodology for developing predictive models to classify the antioxidant activity of phenolic compounds, leveraging Decision Tree-based algorithms and Conceptual Density Functional Theory (CDFT) descriptors. The resulting models achieved high accuracy and full explainability through explicit, interpretable if–then rules derived from molecular features.
6.4 Mixed Explanation
@@ 7 Explaining Sensory Characteristics
Sensory characteristics are extremely important for quality control, leading to the widespread use of sensors designed to mimic human senses. Among these, spectral devices—commonly used and established in the industry—offer rich information, suggesting potential applications for XAI techniques. However, it was observed that most studies focus on pictorial data and visual explanations, with only two works to date addressing spectral data for explainability. In Table 4 we summarize the studies surveyed in the present section.
Table 4: Summary of the works introducing applying XAI for sensory characteristics surveyed in Section 7, according to their data type and explanation type (labeled as “Expl. type”).
@@
[195, 196, 197] Tabular Numerical
7.1 Visual Explanation
@@
Beyond fruit, cereal integrity has been explored using XAI methods. [187] applied Grad-CAM in an EfficientNet-B3-DAN model to detect rice germ integrity, confirming the model’s focus on relevant features. [185] addressed crop yield estimation by developing an Inception-ResNet-based regression model for leaf counting, handling occlusions in monocots. Grad-CAM analysis confirmed its focus on leaf tips, validating effectiveness across sorghum and maize datasets. [191] enhanced crop classification by employing a MobileNetV2 model validated with Grad-CAM to assess the visual standard quality of tomatoes, classifying them as damaged, old, ripe, and unripe. The theme of product freshness is also explored in [190], [189], [192]. [190] introduced a DL-based model to classify meat freshness into fresh, half-fresh, and spoiled categories, incorporating Grad-CAM++ to support transparent decision-making. [189] presented an InceptionV3 model combined with LIME for efficient and transparent classification of chicken meat freshness, which, when integrated with a robotic arm, enhances automation and food safety in poultry processing. [192] utilized CNN-based models to predict the quality of seabream—categorized as fresh, moderate, or spoiled—based on eye and gill images taken under refrigerated conditions, incorporating LIME and Grad-CAM for model interpretability.
@@
Unlike the previous study, two studies used XAI techniques to analyze spectral data to address sensory characteristic problems. [193] developed a CNN model to classify beef freshness using myoglobin data and reflectance spectra, achieving high F1-scores. Grad-CAM highlighted key wavelength regions, confirming myoglobin’s importance in freshness classification. The method demonstrated robustness against environmental factors, indicating strong industrial potential. Similarly, [194] used surface-enhanced Raman spectroscopy and a CNN-based model, the Dual-Branch Wide Kernel Network, to classify bacterial signals.
7.2 Numerical Explanation
@@
Three works address Sensory Characteristics with the goal of explaining model outputs through numeric explanations, despite their differing approaches and applications based on tabular data. [195] focused on predicting boar taint, an undesirable taste and odor found in the meat of male pigs. Using CatBoost, a tree-based ensemble model, the authors achieved peak performance. SHAP analysis identified key factors correlated with boar taint, including feed type, ventilation system, pharmaceutical treatment, and lairage waiting time.[196] developed DL models to classify sweet, bitter, and umami molecules, employing a DNN with molecular descriptors and a graph NN, achieving similar accuracies. SHAP analysis was applied to interpret DNN predictions, revealing key molecular binding properties. [197] developed an ML-based method using various regression models, including XGBoost and Random Forest, alongside Feature Importance analysis, to predict aroma partitioning in dairy matrices and support food reformulation efforts.
@@ 8 Explaining Sustainability and Healthiness
Table 5: Summary of the works introducing applying XAI for sustainability and healthiness surveyed in Section 8, according to their data type and explanation type (labelled as “Expl. type”).
@@
[198, 199, 200, 201, 202, 203] Pictorial Visual
[204, 205, 206, 207, 208, 209, 210, 211] Tabular Numerical
[34] Time series Numerical
@@
8.1 Visual Explanation
The works in this section used Grad-CAM as an XAI technique, confirming its widespread application in explaining solutions to sustainability and healthiness problems using pictorial data. [198] used a CNN combined with a feature-based cascade classifier to achieve 83838383% accuracy in pig face recognition. They employed Grad-CAM to verify that the model focuses on key facial features, offering a cost-effective alternative for animal identification in intensive farming. The paper contributes to the field of animal identification, improving welfare and non-invasive animal management practices. [201] utilized a CNN model, AlexNet, with UAV-based RGB imagery to predict forage biomass, achieving a Mean Absolute Error of 12.9812.9812.9812.98%, with Grad-CAM confirming that the model accurately identified relevant regions for biomass prediction. [200] proposed a CNN model to detect rice phenology stages using smartphone images, reaching 91.3091.3091.3091.30% accuracy. Grad-CAM showed that the model effectively recognized developmental stages, demonstrating the potential of using low-cost tools for real-time agricultural monitoring. [199] introduced MSANet, a model combining multiscale attention and CNN layers for fruit recognition. Grad-CAM was used to interpret the model’s decisions, ensuring effective feature identification for robust fruit classification across applications. This work advances waste reduction through automated fruit detection, promoting environmental sustainability. Similarly, [203] applied a Vision Transformer (ViT) model for plant seedling classification and used attention heatmaps to provide insights into the model’s decision-making process. Lastly, [202] developed a CNN-based system as an automated method for evaluating the precision of agricultural sprayers by detecting spray deposits, eliminating the need for manual tracers or water-sensitive papers. The study also employed an XAI pipeline—specifically Grad-CAM and Grad-CAM++—to interpret the CNN’s decision-making process, revealing key spatial filtering methods used for classification.
@@ 8.2 Numerical Explanations
The studies explored the application of ML and XAI techniques in health, food, and agriculture using tabular data. [205] employed an RF model, using SHAP values to assess the impact of phenol-enriched olive oils on cardiometabolic health in hypercholesterolemic individuals. The study found that phenol-enriched oils significantly reduced serum metabolites associated with cardiovascular risk, indicating their potential as a treatment for cardiometabolic diseases. [204] predicted Oral Food Challenge (OFC) outcomes for diagnosing food allergies, with Random Forest and Learning Using Concave and Convex Kernels models achieving high accuracy in identifying egg, peanut, and milk allergies. SHAP analysis highlighted key clinical factors, such as Immunoglobulin E levels, as important predictors of OFC outcomes. [206] combined genomic and environmental data to predict wheat yield using advanced DL frameworks.
@@
DeepLift [212] analysis revealed that environmental factors were more influential than genetics, highlighting the importance of integrating both data types for crop variety development. [211] integrated ML and DL models—including SVM, RF, and neural networks—with LIME and SHAP to provide a transparent and efficient solution for crop yield prediction, focusing on automating agricultural processes and promoting sustainability. [209] exploited ML-based techniques integrated with LIME and SHAP to predict cattle behavior using sensor data collected from eighteen cows via accelerometers and pressure sensors, classifying behaviors into Other behavior, Ruminating, and Drinking/Eating. [207] applied an RF model to predict almond shelling fraction using genotype data, with SHAP analysis offering insights into the genetic markers influencing shelling fraction, thereby supporting informed breeding strategies. [208] proposed the use of a sensing agricultural robot that collects data such as temperature, humidity, and UV index to automatically forecast mulberry plant diseases by monitoring environmental conditions over time, leveraging LightGBM for prediction and SHAP for interpretability. Finally, [210] introduced a real-time irrigation management system for paddy fields, utilizing a hybrid and ensemble feature extraction approach (HyEn-X) combined with a Federated Learning-based framework, enabling decentralized learning for localized decision-making while preserving data privacy; SHAP was employed to enhance model interpretability.
@@
Considering time series, [34] proposed several ML models to predict individual pig growth trajectories from group-level weight data, reducing reliance on traditional, costly Radio Frequency Identification tracking. The Random Forest model performed best, with an average Root Mean Square Error of $2.262.26$ kg per pig. SHAP analysis highlighted weight and time differences as key predictors, supporting ML as a cost-effective alternative for growth estimation.
8.3 Rule-based Explanation
@@
8.4 Mixed Explanation
@@ 9 Comparison and Insights
Figure 4: Distribution of papers surveyed in the present work per publication year and data type. Most of the articles were published from 2020 onwards, illustrating how the application of XAI to food quality topics is a rather new and evolving discipline.
@@
Figure 5: Distribution of papers surveyed in the present work per topic and explanation type.
@@
Figure 6: Alluvial plot showing the distribution of works surveyed per topic, data type, and explanation type, as introduced in Section 2. The numbers next to the labels indicate the number of works in the specific category. While the distribution across topics is rather uniform, most of the works we survey concentrate on pictorial data and visual explanations, while a smaller portion of research deals with tabular data and numerical explanations.
@@
Figure 7: Pie chart illustrating the distribution of papers in this survey that utilize global XAI techniques versus local XAI techniques (a), and another pie chart depicting the percentages of papers that apply model-agnostic XAI techniques compared to model-specific XAI techniques (b). In both charts, the Both sector indicates that there are works employing multiple techniques of different types.

### v2

@@ 5.3 Mixed Explanation for the Authenticity and Traceability topic
Time series data.
@@
Spectral data.
Using spectral data, [85] investigated 1H NMR spectra to determine the geographical origins of Asian red pepper powders, employing ML, SVM, and CNN models with dimensionality reduction techniques. Grad-CAM and SHAP provided insights into the decision-making processes, highlighting metabolite distribution variations as key classification factors. This study demonstrated the potential of these models for broader applications in food authenticity verification.
@@ 6 Explaining Nutritional Value
Table 5: Summary of the works introducing applying XAI for the Nutritional Value topic surveyed in Section 6, according to their data type and explanation type (labelled as “Expl. type”).
@@
6.1 Visual Explanation for the Nutritional Value topic
Pictorial data.
@@ 6.1 Visual Explanation for the Nutritional Value topic
6.2 Numerical Explanation for the Nutritional Value topic
Tabular data.
Tabular data was discussed in [52], [76], [60], and [105], which apply ML techniques to various food-related challenges. [52] used XGBoost to estimate added sugar content in foods, with SHAP enhancing model transparency for consumer awareness in regions without mandatory labeling. [76] developed the Flavonoid Astringency Prediction Database, employing ML models like RF to explore the relationship between molecular structures and flavor properties. Similarly, [60] applied ML to differentiate pepper spices during storage, using SHAP to identify key organic compounds. [105] developed an XGBoost-based model for predicting drug-food interactions using molecular fingerprint similarities, with SHAP providing insights into influential features relevant to clinical applications and dietary planning. [58] proposed a graph-based ML approach to predict the outcomes of formulation trials, aiming to reduce laboratory experiments, material waste, and development time in food design. To enhance interpretability, they applied GNNExplainer [262], a global explanation method tailored for Graph NNs.
Spectral data.
Spectral data was explored in [101]. The authors employed visible NIR point spectroscopy to estimate sugar content in grape varieties at different maturity stages. Regression ML algorithms and a CNN were applied, with XAI techniques such as Variable Importance in Projection and Gini Importance validating the models and identifying key spectral features.
6.3 Rule-based Explanation for the Nutritional Value topic
Pictorial data.
[193] exploited pictorial data to propose a similarity score based on user community preferences, enhancing recommendation quality. The rule-based explainability method assigned each image to an appropriate food diet based on user profiles, supporting personalized dietary recommendations.
Tabular data.
[77] presented a novel no-code methodology for developing predictive models to classify the antioxidant activity of phenolic compounds, leveraging Decision Tree-based algorithms and Conceptual Density Functional Theory (CDFT) descriptors. The resulting models achieved high accuracy and full explainability through explicit, interpretable if–then rules derived from molecular features.
6.4 Mixed Explanation for the Nutritional Value topic
Tabular data.
@@ 7 Explaining Sensory Characteristics
Sensory characteristics are extremely important for quality control, leading to the widespread use of sensors designed to mimic human senses. Among these, spectral devices—commonly used and established in the industry—offer rich information, suggesting potential applications for XAI techniques. However, most studies focus on pictorial data and visual explanations, with only two works to date addressing spectral data for explainability. In Table 6 we summarize the studies surveyed in the present section.
Table 6: Summary of the works introducing applying XAI for the Sensory Characteristics topic surveyed in Section 7, according to their data type and explanation type (labeled as “Expl. type”).
@@
[134, 61, 9, 176, 261, 260] Tabular Numerical
[212] Tabular Rule-based
7.1 Visual Explanation for the Sensory Characteristics topic
Pictorial data.
@@ 7.1 Visual Explanation for the Sensory Characteristics topic
Beyond fruit, cereal integrity has been explored using XAI methods. [116] applied Grad-CAM in an EfficientNet-B3-DAN model to detect rice germ integrity, confirming the model’s focus on relevant features. [254] addressed crop yield estimation by developing an Inception-ResNet-based regression model for leaf counting, handling occlusions in monocots. Grad-CAM analysis confirmed its focus on leaf tips, validating effectiveness across sorghum and maize datasets. [205] enhanced crop classification by employing a MobileNetV2 model validated with Grad-CAM to assess the visual standard quality of tomatoes, classifying them as damaged, old, ripe, and unripe.
The theme of product freshness is also explored in [224], [82], [73]. [224] introduced a DL-based model to classify meat freshness into fresh, half-fresh, and spoiled categories, incorporating Grad-CAM++ to support transparent decision-making. [82] presented an InceptionV3 model combined with LIME for efficient and transparent classification of chicken meat freshness, which, when integrated with a robotic arm, enhances automation and food safety in poultry processing. [73] utilized CNN-based models to predict the quality of seabream—categorized as fresh, moderate, or spoiled—based on eye and gill images taken under refrigerated conditions, incorporating LIME and Grad-CAM for model interpretability.
A different study, which employed similar techniques, specifically a ResNet with Grad-CAM to explain the results, was proposed by [55]. The authors introduced a framework to monitor the storage stability of beeswax oleogel using computer vision based on DL trained on polarized-light microscopy images of the oleogel.
Spectral data.
@@
Two studies used XAI techniques to analyze spectral data to address sensory characteristic problems. [213] developed a CNN model to classify beef freshness using myoglobin data and reflectance spectra, achieving high F1-scores. Grad-CAM highlighted key wavelength regions, confirming myoglobin’s importance in freshness classification. The method demonstrated robustness against environmental factors, indicating strong industrial potential. Similarly, [190] used surface-enhanced Raman spectroscopy and a CNN-based model, the Dual-Branch Wide Kernel Network, to classify bacterial signals.
7.2 Numerical Explanation for the Sensory Characteristics topic
Tabular data.
@@ 7.2 Numerical Explanation for the Sensory Characteristics topic
Several works address Sensory Characteristics topic with the goal of explaining model outputs through numeric explanations, despite their differing approaches and applications based on tabular data. [134] focused on predicting boar taint, an undesirable taste and odor found in the meat of male pigs. Using CatBoost, a tree-based ensemble model, the authors achieved peak performance. SHAP analysis identified key factors correlated with boar taint, including feed type, ventilation system, pharmaceutical treatment, and lairage waiting time.[61] developed DL models to classify sweet, bitter, and umami molecules, employing a DNN with molecular descriptors and a Graph NN, achieving similar accuracies. SHAP analysis was applied to interpret DNN predictions, revealing key molecular binding properties. [9] developed an ML-based method using various regression models, including XGBoost and RF, alongside Feature Importance analysis, to predict aroma partitioning in dairy matrices and support food reformulation efforts. [261] and [260] proposed two similar applications of an AI framework for modelling and interpreting the rheological behaviour of plant-protein systems under varying processing and thermal conditions. More specifically, they apply regression models for predicting shear stress, viscosity, and temperature-dependent viscoelastic properties of sesame protein isolate gels from processing, thermal, and compositional variables. Moreover, SHAP is applied to explain the outcomes. [176] introduced an ML-based method for predicting dry matter content in Emmental cheese, again using SHAP to explain the obtained results.
7.3 Rule-based Explanation for the Sensory Characteristics topic
Tabular data.
There is a single example of applying XAI to Sensory Characteristic tasks, and it is based on tabular data. [212] developed an intrinsically interpretable model to predict crab weight using easily measurable physical features. The proposed model was based on genetic programming, a class of evolutionary algorithms that automatically evolve computer programs (often represented as trees) to solve a task using operations analogous to biological evolution, such as selection, crossover, and mutation [177]. The output of the model is a symbolic regression formula that relates crab weight to its characteristics in a transparent and readable way.
@@ 8 Explaining Sustainability and Healthiness
Table 7: Summary of the works introducing applying XAI for the Sustainability and Healthiness topic surveyed in Section 8, according to their data type and explanation type (labelled as “Expl. type”).
@@
[138, 152, 79, 35, 192, 68, 266] Pictorial Visual
[67, 99, 167, 231, 89, 219, 98, 114, 93] Tabular Numerical
[227, 37] Time series Numerical
@@
8.1 Visual Explanation for the Sustainability and Healthiness topic
Pictorial data.
The works in this section used Grad-CAM as an XAI technique, confirming its widespread application in explaining solutions to Sustainability and Healthiness problems using pictorial data. [138] used a CNN combined with a feature-based cascade classifier in pig face recognition. They employed Grad-CAM to verify that the model focuses on key facial features, offering a cost-effective alternative for animal identification in intensive farming. The paper contributes to the field of animal identification, improving welfare and non-invasive animal management practices. [35] utilized a CNN model, AlexNet, with UAV-based RGB imagery to predict forage biomass, with Grad-CAM confirming that the model accurately identified relevant regions for biomass prediction. [79] proposed a CNN model to detect rice phenology stages using smartphone images. Grad-CAM showed that the model effectively recognized developmental stages, demonstrating the potential of using low-cost tools for real-time agricultural monitoring. [152] introduced MSANet, a model combining multiscale attention and CNN layers for fruit recognition. Grad-CAM was used to interpret the model’s decisions, ensuring effective feature identification for robust fruit classification across applications. This work advances waste reduction through automated fruit detection, promoting environmental sustainability. [266] created a modular DL-based framework, called AMULET, that predicts plant traits from RGB images. The framework was tested on potato plants and on Arabidopsis thaliana, but could be extended to other plant species by predicting plant appearance. Two gradient-based XAI techniques, TorchGrad and Grad-CAM, were applied to explain the results. Similarly, [68] applied a Vision Transformer (ViT) model for plant seedling classification and used attention heatmaps to provide insights into the model’s decision-making process. Lastly, [192] developed a CNN-based system as an automated method for evaluating the precision of agricultural sprayers by detecting spray deposits, eliminating the need for manual tracers or water-sensitive papers. The study also employed an XAI pipeline—specifically Grad-CAM and Grad-CAM++—to interpret the CNN’s decision-making process, revealing key spatial filtering methods used for classification.
@@ 8.2 Numerical Explanations
Tabular data.
@@
The studies explored the application of ML and XAI techniques in health, food, and agriculture using tabular data. In this context, the works primarily address sustainable cultivation and livestock management practices, with particular attention to animal behavior. [98] integrated ML and DL models—including SVM, RF, and NNs—with LIME and SHAP to provide a transparent and efficient solution for crop yield prediction, focusing on automating agricultural processes and promoting sustainability. [99] combined genomic and environmental data to predict wheat yield using advanced DL frameworks. DeepLift [216] analysis revealed that environmental factors were more influential than genetics, highlighting the importance of integrating both data types for crop variety development. [231] proposed the use of a sensing agricultural robot that collects data such as temperature, humidity, and UV index to automatically forecast mulberry plant diseases by monitoring environmental conditions over time, leveraging LightGBM for prediction and SHAP for interpretability. [194] proposed a method to determine irrigation requirements for spinach using soil and climate data. The model was based on a stacked NN architecture, specifically a CNN and a Long Short-Term Memory network, to predict irrigation demands. Similarly, [219] introduced a real-time irrigation management system for paddy fields, utilizing a hybrid and ensemble feature extraction approach, called HyEn-X, combined with a Federated Learning-based framework, enabling decentralized learning for localized decision-making while preserving data privacy; SHAP was employed to enhance model interpretability.
[89] exploited ML-based techniques integrated with LIME and SHAP to predict cattle behavior using sensor data collected from eighteen cows via accelerometers and pressure sensors, classifying behaviors into Other behavior, Ruminating, and Drinking/Eating. [93] proposed a DL system that predicted the risk of ducks dying on arrival during transport in trucks. SHAP played an essential role in understanding the contribution of the features learned by the model.
[67] focuses on human health and potential risks. The authors employed an RF model, using SHAP values to assess the impact of phenol-enriched olive oils on cardiometabolic health in hypercholesterolemic individuals. The study found that phenol-enriched oils significantly reduced serum metabolites associated with cardiovascular risk, indicating their potential as a treatment for cardiometabolic diseases.
A couple of studies focus on plant phenotypes. [167] applied an RF model to predict almond shelling fraction using genotype data, with SHAP analysis providing insights into the genetic markers influencing shelling fraction and thereby supporting informed breeding strategies. citenr193 developed a system for field-scale potato phenotyping by leveraging UAV-based multispectral imagery and machine learning. The authors engineered novel features from the spectral data and used XGBoost and RF for prediction. SHAP confirmed that the proposed approach provides a non-invasive and highly accurate method for estimating potato plant traits.
Time series data.
@@
Considering time series, [227] proposed several ML models to predict individual pig growth trajectories from group-level weight data, reducing reliance on traditional, costly Radio Frequency Identification tracking. The RF model performed best, with an average Root Mean Square Error of $2.26$ kg per pig. SHAP analysis highlighted weight and time differences as key predictors, supporting ML as a cost-effective alternative for growth estimation. [37] introduced a system for monitoring equine behavior based on wearable sensors and ML. In particular, they proposed using RF to predict animal behavior (e.g., walking, grazing, galloping, head shaking) using data collected from a neck-mounted accelerometer, gyroscope, and magnetometer. SHAP was used to explain the obtained results.
8.3 Rule-based Explanation for the Sustainability and Healthiness topic
Tabular data.
@@ 8.3 Rule-based Explanation for the Sustainability and Healthiness topic
Time series data.
[234] highlighted the need to monitor low-cost, automated, and interpretable irrigation systems using time series data. To address this, they proposed a new system called Vital, which integrates IoT sensors, a data management platform, and a fuzzy rule-based decision support system to automate irrigation. The system was evaluated through pilot cases and effectively automated the irrigation process, monitoring, and managing open-field installations that provided water.
8.4 Mixed Explanation for the Sustainability and Healthiness topic
Tabular data.
@@ 9 Comparison and Insights
Figure 6: Distribution of surveyed papers by publication year and data type. The sharp increase from 2020 onwards reflects the growing interest in applying XAI techniques to food quality, with pictorial data dominating early research and tabular data gaining traction in recent years. Time series and spectral data remain underexplored despite their widespread use in physicochemical analysis.
@@
Figure 7: Distribution of surveyed papers by food quality topic and explanation type. Food Security emerges as the dominant application area, followed by Authenticity and Traceability and Nutritional Value. Sensory Characteristics and Sustainability and Healthiness remain underexplored, and a consistent gap is observed in the use of XAI for spectral and time series data across all topics.
@@
Figure 8: Distribution of works surveyed. While the distribution across topics is rather uniform, most of the works we survey concentrate on pictorial data and visual explanations, while a smaller portion of research deals with tabular data and numerical explanations.
@@
Figure 9: Distribution of surveyed papers that utilize global XAI techniques versus local XAI techniques (a), and the percentages of papers that apply model-agnostic XAI techniques compared to model-specific XAI techniques (b). In both charts, the Both sector indicates that there are works employing multiple techniques of different types. Observing the first graph, it becomes evident that most of the studies reviewed employ local XAI techniques, such as the widely recognized SHAP and Grad-CAM. In contrast, global techniques appear to be less commonly used. The second graph confirms this observation, since SHAP is model-agnostic, while Grad-CAM is model-specific, highlighting an even distribution of the two most popular techniques across the various works.
@@
9.1 Critical evaluation of XAI methods
While XAI techniques can enhance transparency and support model validation, their outputs should not be interpreted as ground truth. In food engineering, where decisions may affect safety, compliance, and product quality, the reliability of explanations becomes a central concern. In the following section, we examine the principal limitations and practical implications of the most widely adopted explainability techniques, as identified through the systematic analysis of the surveyed literature. These limitations should therefore not be regarded as inherent properties of the techniques themselves or as universally valid claims, but rather as recurring patterns that emerged from the examination of the reviewed works.
Reliability and stability of model-agnostic explainers (LIME/SHAP).
Model-agnostic explainers such as LIME and SHAP are widely adopted because they can be applied to different black-box models and data types. However, both methods are susceptible to instability and sensitivity with respect to design choices [160]. For LIME, explanations depend on how the local neighborhood is sampled, including the number of perturbed instances, the kernel width, and the distance metric, as well as on the choice of perturbation operator. Small variations in these settings can yield different surrogate models and, consequently, divergent feature attributions. This issue is particularly problematic in high-dimensional settings such as spectral vectors or image inputs, where perturbations can generate unrealistic samples that lie outside the data manifold, thereby undermining the quality of the local surrogate.
Similarly, SHAP explanations are sensitive to the background or reference distribution used to simulate feature removal, as well as to the assumptions made regarding feature dependence. In food datasets, predictors are frequently correlated, as is the case with physicochemical measurements, spectral wavelengths, and environmental covariates. Independence-based approximations can therefore distribute attribution credit across correlated features in a manner that does not accurately reflect the model’s actual decision logic. Furthermore, SHAP values may shift when the baseline is altered, for instance, when substituting the training-set mean with a task-specific reference point, which can undermine interpretability for domain practitioners.
When employing LIME or SHAP in food quality applications, it is therefore important to explicitly report the perturbation strategy and baseline selection, and to treat the resulting attributions as hypotheses to be validated rather than as definitive explanations.
Noise and sensitivity in saliency-based visual explanations.
Saliency and CAM-based techniques, including CAM, Grad-CAM, and their variants, are widely used for visual data because they produce heatmaps that are intuitive for human stakeholders. Nevertheless, saliency maps can be noisy and sensitive to model parameters, input preprocessing choices, and small input perturbations.
In practice, heatmaps may highlight broad or diffuse regions and spurious patterns such as backgrounds, lighting artifacts, packaging elements, or imaging device signatures, particularly when models exploit shortcut cues during training. This risk is especially relevant in food inspection scenarios where acquisition conditions vary across illumination, moisture levels, occlusion, and camera angle, and where visually salient regions do not necessarily correspond to causally meaningful factors underlying the quality attribute of interest. Additionally, CAM-based methods typically provide coarse spatial localization and should not be interpreted as precise segmentation of defects or contaminated areas.
As a mitigation strategy, studies employing visual explanation methods can complement heatmap analysis with sanity checks such as randomization tests, sensitivity analysis under mild input perturbations, and systematic comparisons across multiple explainers to assess the consistency and reliability of the resulting attributions.
Recommendations and limitations for the effective application of XAI techniques.
The choice of XAI method should be aligned with the data modality, the model architecture being employed, and the informational needs of the relevant stakeholders. For instance, the present survey reveals that when NNs are applied to pictorial data, the selection of an XAI technique falls almost exclusively on Grad-CAM, despite the fact that other established methods perform comparably well in this setting. A more informed use of the available XAI techniques, grounded in a clear understanding of what each method actually computes, can help overcome the substantial difficulties that commonly limit their effective adoption.
One such difficulty is the gap between a plausible explanation, that is, one that appears convincing to a human observer, and an explanation that genuinely reflects the model’s true decision process. In some cases, an explanation may seem counterintuitive or unconvincing while nonetheless accurately capturing the model’s learned behavior. In others, an explanation may appear well-aligned with user expectations while failing to correspond to what the model has actually learned [12]. In food engineering, this gap is amplified by domain complexity and the practical urgency of operational decisions.
We therefore recommend treating XAI as a diagnostic layer rather than a source of ground-truth justifications. Explanations should be triangulated against domain knowledge, controlled experiments, and independent measurements where feasible. Furthermore, combining explanations from multiple methodological families, such as SHAP and PDP for tabular data, or Grad-CAM paired with a perturbation-based method for image data, provides a more robust characterization of model behavior and reduces the risk of over-interpreting artifacts produced by any single XAI approach.
Finally, we recommend extending the application of XAI techniques to every stage of the pipeline in which an ML or AI model is employed. In practice, preprocessing and data cleaning steps frequently rely on models that themselves operate as black boxes, yet their outputs propagate through the entire workflow and can substantially influence the final result. We therefore advocate for the consistent application of explainability methods to these upstream stages as well, so as to ensure transparency and reliability throughout the full analytical pipeline rather than solely at the prediction stage [36].
Recommendations for reporting and validation.
To improve transparency and reproducibility, we recommend that food quality studies employing explainability methods explicitly report the explainer configuration, including the perturbation operator, kernel width, and baseline choices, sensitivity analyses assessing the stability of explanations under resampling or mild input perturbations, and quantitative metrics that justify both the appropriateness of the chosen technique and the validity of the resulting explanations, such as faithfulness-oriented evaluations through removal-based assessments and sanity checks for saliency maps, along the lines of the evaluation framework proposed by [160].